An unknown compound has the formula CxHyOz. When 0,1523 g of the compound is burned in oxygen, 0,3718 g of CO2 and 0,1522 g of H2O is produced. 

(a) What is the empirical formula of the unknown compound? 

(b) If the molar mass of the unknown compound is 72,1 g/mol, what is its molecular formula?  



1
Expert's answer
2021-04-21T04:05:08-0400

"Moles (O_2) =\frac{0.1523}{16}=9.5×10^{-3}"


"Moles (Co_2) =\frac{0.3718}{44}=8.5×10^{-3}"


"Moles (O_2) =\frac{0.1522}{18}=8.5×10^{-3}"



Hence

"\frac{9.5×10^{-3}}{8.5×10^{-3}}=1.1"


"\frac{8.5×10^{-3}}{8.5×10^{-3}}=1"


"\frac{8.5×10^{-3}}{8.5×10^{-3}}=1"


Emperical formula = "CHO_2"


"Moles =\frac{ 0.1523}{72.1}=2.1×10^{-3}"


Molecular formula "= C_2H_2O_4"